OpenStudy (anonymous):

Write an equation of the line in point slope form that passes through the given points. Points: (-15,7) (5,-3)

OpenStudy (texaschic101):

first we need to find the slope by using the slope formula slope(m) = (y2 - y1) / (x2 - x1) (-15,7)....x1 = -15 and y1 = 7 (5,-3)....x2 = 5 and y2 = -3 now sub slope(m) = (-3-7) / (5 - (-15) can you finish this and find the slope ?

OpenStudy (texaschic101):

whenever you are ready to participate, I am here to help :)

OpenStudy (icecreamyummy):

Y=-.5+-.5